OpenStudy (anonymous):

Hey guys, Please help me with this question. Thanks! The tortoise and the hare are running a 2000 m race. The tortoise travels 20 m every minute. The tortoise proceeds very slowly, never changing its speed.The hare travels 2000m in the first minute but, in each minute after that, travels only half of the remaining distance ( in the 2nd minute, the hare travels half of the remaining 1000 m, and so on). a) Write the series for the distance the tortoise travels for the first few minutes of the race, where S(n) represents the distance the tortoise travels in that minute. b) Write the series fo

OpenStudy (anonymous):

I don't get it, if the hare run at 2000 m first minutes, wouldn't it finished the race?

OpenStudy (anonymous):

each minute the hare is going half the distance . so after 200o metres he is going 100 metres next minute and then 500

OpenStudy (anonymous):

but the tortoise is going at a constant spped of 20 metres per minute. i have to find bothe sequencces for the tortoise and the hare and then calculate who is going to win the race. please help me with this

OpenStudy (anonymous):

Oh, you mean 200, 100, 50

OpenStudy (anonymous):

the hare 2000, 1000, 500, 250, 125....... the tortoise is going at constant speed 20, 20, 20 ,20,20

OpenStudy (anonymous):

Hare \[\sum _{n=0}^N 2000\left(\frac{1}{2}\right)^n\]

OpenStudy (anonymous):

Tortoise \[\sum _{n=0}^N 20\]

OpenStudy (anonymous):

so who is going to win the race? i need the S(n) or the total distance covered by the hare and the tortoise to prove who won the race in how many minutes

OpenStudy (anonymous):

I know this go against the story but hare is

OpenStudy (anonymous):

so buddy in which minute does the hare reach the finishing line???

OpenStudy (anonymous):

since the race is 2000 m and hare ran 2000 in first minutes, hare reached the finished line in first minute